OpenStudy (anonymous):

write the eqaution of the line in standard form? Has a slope of 8 and passes through the point (2, -3)??

OpenStudy (anonymous):

(y-(-3))=8*(x-2) y+3=8*(x-2)

OpenStudy (anonymous):

y=8x-19